# Cosmic-Ray Feedback from Supernovae in a Stratified Interstellar Medium

### v1

@@ 4.1 Cosmic-Ray Calorimeter Problem
The steady state structure and change in cosmic-ray transport we see in the CRInj simulation could provide a new solution to the problem of large $\gamma$-ray luminosities in global galaxy simulations. This problem was examined in depth with the FIRE simulations in Chan et al. (2019). They found that the only way to match the observed $\gamma$-ray luminosities of lower stellar surface density galaxies was to increase the cosmic-ray diffusion coefficient. But the fundamental problem was one of calorimetry. Observational estimates have shown that smaller (lower $\Sigma_{*}$) galaxies are not cosmic-ray calorimeters, requiring $F_{\mathrm{cal}}<1$ (Lacki et al., 2011). But all of the streaming and Milky-Way diffusion coefficient simulations in Chan et al. (2019) were at or above the calorimetric limit. The “cosmic-ray calorimeter problem” is that simulated galaxies end up being cosmic-ray calorimeters. However, we calculated the $F_{\mathrm{cal}}$ parameter from Lacki et al. (2011) and found that our simulation got close to their observational estimates for starbursting galaxies (see top panel of Figure 17).
@@
Our $F_{\mathrm{cal}}$ parameter gets close to the observed values without increased diffusion. Our cosmic rays are transported diffusively along the local magnetic field direction, like some of the simulations in Chan et al. (2019). However, combining this anisotropic transport with a higher resolution in the diffuse gas allowed the cosmic rays to escape along vertically oriented magnetic field lines. We suspect resolution is the reason this has not been noted before. Since we resolve the entire simulation at $\Delta x=10\,\mathrm{pc}$, we resolve more of the magnetic field structure outside the midplane. For comparison with finite mass codes, a gas mass resolution of $m_{g}\approx 0.025M_{\odot}$ would be necessary to resolve $n=10^{-3}\mathrm{cm}^{-3}$ gas at $\Delta x=10\,\mathrm{pc}$. This mass resolution is much smaller than the $m_{g}\approx 10^{3}M_{\odot}$ used in galaxy simulations like Chan et al. (2019).
@@
However, the $F_{\mathrm{cal}}$ parameter rises again once we get numerical reconnection which reduces the amount of vertical field. The numerical magnetic reconnection occurs where the Parker instability plumes run into each other. While there is still some vertical flux which survives (see Figure 13), the reconnection decreases the total magnetic energy density (see Figure 3). It is necessary to study this process in higher resolution to determine what happens when the effects of numerical reconnection are minimized.
@@
We suggest others also reconsider the resolution of diffuse gas outside the midplane of the galaxy when examining the cosmic-ray calorimeter problem. While the gas dynamics do not require a high resolution, it might be necessary to produce a realistic magnetic field structure. Naively, we would expect magnetic field lines could stretch from the midplane, out and up into the circumgalactic medium. Without higher resolution in the diffuse gas, these vertical field lines numerically reconnect and the region becomes disconnected from the midplane. But if the vertical field lines were still there, connecting the midplane to the diffuse gas, then they would provide an escape highway for cosmic rays. These hidden highways could be the true solution to the calorimeter problem, instead of an increased cosmic-ray diffusion coefficient.
@@ 4.2 Cosmic-Ray Transport and Stability of the Disk-Halo Interface
In our simulations, the cosmic rays are predominantly diffusive at high densities but are transported by streaming at low densities (see top row of Figure 14). This results from the density dependence of the Alfven speed. At lower densities, the Alfven speed gets larger, allowing the streaming transport to be larger than diffusive transport. Because of the steady stratification, this density dependence is also a vertical position dependence. In our simulations, we see cosmic rays are streaming outside of the midplane. This change in transport matches the recent results from other groups who identify the same switch in cosmic-ray transport outside the midplane (Armillotta et al., 2024; Thomas et al., 2024).
@@
A change in cosmic-ray transport also means a change in feedback. When cosmic rays are streaming at the Alfven speed, they heat the gas (see middle and bottom rows of Figure 14) by scattering off of Alfven waves they generate themselves. Additionally, the streaming process is an advective process instead of a diffusive process. Instead of smoothing out cosmic-ray pressure gradients like diffusion, any steep cosmic-ray pressure gradients are advected at the Alfven speed. So the switch to streaming means larger cosmic-ray pressure gradients, creating a larger force on the thermal gas, which can drive outflows (see Figure 10). These impacts of streaming have been discussed extensively in other works (e.g. Tsung et al. 2023).
@@
The $P_{c}$-$n$ diagrams in the top row of Figure 14 point to a new idea regarding cosmic-ray feedback at the disk-halo interface. When cosmic-ray transport changes, it also changes the effective polytropic index of cosmic rays. A change in polytropic index changes the stability of the system (note the appearance of $\gamma_{c}$ in Equation 21). Our simulations show an increase in polytropic index with height. The effective polytropic index $\gamma_{c}$ shows up on the right hand side of Equation 21, meaning an increase has a stabilizing effect. So, the change in cosmic-ray transport with $z$ could naturally cause a fountain flow: gas parcels moving upward because they are unstable before eventually reaching a position where their compressibility increases and they become stable to perturbations.
@@
We developed this interpretation after reading Hosking et al. (2024), which illustrated how a plasma’s compressibility can impact the structure of metastable, stratified MHD atmospheres. A detailed analysis on the stability of a CR+MHD atmosphere is out of the scope of this paper. Instead, we just share the idea here and will examine it in detail in a future work (Habegger, Hosking & Zweibel, in prep.).
@@ 5 Conclusion
The simulations we have run, CRInj and NoCRInj, have illustrated the importance of cosmic-ray injections in driving outflows and structuring the ISM. Here, we list some key results from this study of cosmic-ray feedback:
@@
    The Parker instability drives overturning of gas which turns our initial hydrostatic equilibrium into a more realistic two-scale-height vertical gas density structure. This overturning and saturation happens on a shorter timescale than previously predicted: $\sim 150\mathrm{Myr}$ instead of $\sim 500\mathrm{Myr}$ (Heintz et al., 2020)).
@@
    Our simulations produce a steady state after the saturation of the Parker instability, with horizontal magnetic field in the midplane and vertical magnetic field outside the midplane. This magnetic field structure allows for more efficient cosmic-ray escape alongside significant cosmic-ray feedback and a decrease in the calorimetric fraction $F_{\mathrm{cal}}$.
@@
    While the thin gas disk in the steady state is supported by gas pressure and some magnetic pressure, the thick gas disk is supported by non-thermal pressures. The cosmic-ray injections change the outer scale height, inflating the thick disk by a factor of up to $50\%$.
With only two simulations, we cannot be certain of the importance of cosmic-ray injections in all galaxies. In particular, our work does not complete the full stellar feedback loop to show how cosmic-ray injections impact the star formation rate. To complete that loop, the injection rate would at least need to be determined by star formation rates (e.g. like in the galaxy scale simulations of Chan et al. 2019). Because we can not complete the loop, the physical process we highlight of vertical magnetic field lines being cosmic ray escape highways is just a physical explanation of decreased $\gamma$-ray luminosity. The overall effect on star formation rate of magnetic field lines connecting the disk and halo is still unclear. More simulations might illustrate the impact on star formation rate is also the same as in the increased diffusion coefficient case in Chan et al. (2019).
However, we suspect the true picture of cosmic ray feedback on star formation rates will require the increased magnetic field resolution. For example, reconsider Figure 6. In cases where the Parker instability is a dominant dynamical process, the decorrelation of cosmic rays and gas density could significantly impact start formation rates. That decorrlation was caused by field lines being vertical and not undergoing numerical reconnection. As a result, cosmic rays were able to escape the midplane before they could interact with the cold gas.
In the future, we plan to close the feedback loop while exploring the parameter space of these simulations, adjusting initial conditions and quantities listed in Table 1. Additionally, we plan to examine the steady state with a more realistic setup. In addition to tying cold gas formation and injection rate together, we could include other physical processes. Including the effects of galactic rotation (Coriolis force and shearing) would increase the fidelity the magnetic field evolution. A more realistic gravitational field would allow us to determine whether the outflows are just fountain flows or actual winds which escape the galaxy. Additionally, a more accurate heating and cooling function would produce a more realistic multiphase ISM in the simulation.
However, including all these effects will not likely change the fundamental conclusions from this paper. Our goal was to identify how cosmic rays can impact the structure and dynamics of a galactic patch. To do that, we isolated and focused on a few fundamental physical processes. Namely, a resolved magnetic field combined with cosmic-ray injections could help keep a galaxy under the calorimetric limit. The geometry of the Parker instability naturally decorrelated the cosmic rays and dense gas by generating a vertical field alongside cold gas formation. Cosmic rays will also still be efficient at driving an outflow outside of the galactic disk, where their diffusion is no longer dominant because the gas velocities and Alfven speed increase.
The authors would like to thank Sophie Aerdker, Ryan Farber, Ka Wai Ho, David Hosking, Francisco Ley, Nickolas Pingel, Mohan Richter-Addo, Mateusz Ruszkowski, Aaron Tran, Bindesh Tripathi, and Ka Ho Yuen for their insight and helpful conversations which improved this work. The authors greatly appreciate funding from funding from NASA FINESST grant No. 80NSSC22K1749 and NSF grant AST-2007323 which supported this work. Resources supporting this work were provided by the NASA High-End Computing (HEC) Program through the NASA Advanced Supercomputing (NAS) Division at Ames Research Center.

### v2

@@ 4.1 $\gamma$-Ray Bright Galaxy Problem
The steady state structure and change in cosmic-ray transport we see in the CRInj simulation could provide a new solution to the problem of large $\gamma$-ray luminosities in global galaxy simulations. This problem was examined in depth with the FIRE simulations in Chan et al. (2019). They found that the only way to match the observed $\gamma$-ray luminosities of lower stellar surface density galaxies was to increase the cosmic-ray diffusion coefficient beyond the observationally estimated value.
The brightness problem can also be couched as a cosmic-ray calorimeter problem. Observational estimates have shown that smaller (lower $\Sigma_{*}$) galaxies are not cosmic-ray calorimeters, requiring $F_{\mathrm{cal}}<1$ (Lacki et al., 2011). But all of the streaming and Milky-Way diffusion coefficient simulations in Chan et al. (2019) were at or above the calorimetric limit. Therefore, using the observationally expected transport processes, simulated galaxies end up being cosmic-ray calorimeters, producing too many $\gamma$-rays.
We calculated the $F_{\mathrm{cal}}$ parameter from Lacki et al. (2011) and found that our simulation reaches their observational estimates for starbursting galaxies (see top panel of Figure 18), once we account for our initial cosmic ray background.
@@
Our net $F_{\mathrm{cal}}$ parameter gets close to the observed values without cosmic ray diffusion amplified above the observed Milky Way value. Cosmic rays are transported diffusively, and via streaming, along the local magnetic field direction in our simulation, like some of the simulations in Chan et al. (2019). Combining this anisotropic transport with a higher resolution in the diffuse gas allowed the cosmic rays to escape along vertically oriented magnetic field lines. We suspect resolution is the reason this has not been noted before. Since we resolve the entire simulation at $\Delta x=\Delta y=\Delta z=5\,\mathrm{pc}$, we resolve more of the magnetic field structure outside the midplane. For comparison with finite mass codes, a gas mass resolution of $m_{g}\approx 0.129M_{\odot}$ would be necessary to resolve $n=10^{-2}\mathrm{cm}^{-3}$ gas at a resolution of $5\,\mathrm{pc}$. This mass resolution is much smaller than the $m_{g}\approx 10^{3}M_{\odot}$ used in galaxy simulations like Chan et al. (2019).
@@
However, direct calculation of the $F_{\mathrm{cal}}$ parameter (not the net value) suggests our CRInj simulation still has all cosmic ray energy lost to hadronic collisions (see bottom plot of Figure 18). This coincides with a rise in the net $F_{\mathrm{cal}}$ once numerical reconnection occurs, reducing the amount of vertical field. The numerical reconnection occurs where the Parker instability plumes run into each other. While there is still some vertical flux which survives (see Figure 14), the reconnection decreases the total magnetic energy density (see Figure 3). It is necessary to study this process in higher resolution to determine what happens when the effects of numerical reconnection are minimized.
We suggest spatially resolving the diffuse gas outside the midplane of the galaxy is an important ingredient in solving the problem of simulated galaxies being too bright in $\gamma$-ray luminosity. While the dynamics of diffuse gas outside the midplane does not require a high resolution, high resolution might be necessary to produce a realistic magnetic field structure.
@@
Naively, we would expect magnetic field lines should stretch from the midplane, out and up into the galactic halo, and eventually the circumgalactic medium. Without higher resolution in the diffuse gas, these vertical field lines numerically reconnect and the regions become disconnected from the midplane. But if the vertical field lines were still there, connecting the midplane to the diffuse gas, then they would provide an escape highway for cosmic rays. These hidden highways could be the true solution to the calorimeter problem, instead of an increased cosmic-ray diffusion coefficient.
@@ 4.2 Cosmic-Ray Transport and Stability of the Disk-Halo Interface
In our simulations, the cosmic rays are predominantly diffusive at high densities but are transported by streaming at low densities (see top row of Figure 15). At lower densities (in completely ionized gas), the Alfven speed gets larger, allowing the streaming transport to be larger than diffusive transport. Because of the steady stratification, this density dependence is also a vertical position dependence. This change in transport matches the recent results from other groups who identify a similar switch in cosmic-ray transport outside the midplane (Armillotta et al., 2024; Thomas et al., 2024).
@@
A change in cosmic-ray transport also means a change in feedback. When cosmic rays are streaming at the Alfven speed, they heat the gas (see middle and bottom rows of Figure 15) by scattering off of Alfven waves they generate themselves. Additionally, the streaming process is an advective process instead of a diffusive process. Instead of smoothing out cosmic-ray pressure gradients like diffusion, any steep cosmic-ray pressure gradients are advected at the Alfven speed. So the switch to streaming means larger cosmic-ray pressure gradients, creating a larger force on the thermal gas, which can help drive outflows (see Figure 11). These impacts of streaming have been discussed extensively in other works (e.g. Tsung et al. 2023).
@@
The $P_{c}$-$n$ diagrams in the top row of Figure 15 point to a new idea regarding cosmic-ray feedback at the disk-halo interface. When cosmic-ray transport changes, it also changes the effective polytropic index of cosmic rays. A change in polytropic index changes the stability of the system (note the appearance of $\gamma_{c}$ in Equation 21). Our simulations show an increase in polytropic index with height. So, the change in cosmic-ray transport with $z$ could naturally shift the stability of gas parcels moving between the galactic disk and halo.
@@
We developed this interpretation after reading Hosking et al. (2024), which illustrated how a plasma’s compressibility can impact the structure of metastable, stratified MHD atmospheres. A detailed analysis on the stability of a CR+MHD atmosphere is out of the scope of this paper. Instead, we just share the idea here and examine it in detail in a forthcoming paper (Habegger, Hosking & Zweibel, in prep.).
4.3 Caveats
Before concluding, we would like to detail some caveats to this work and the assumptions we made in designing the simulations.
Firstly, the gravitational profile we chose for this equilibrium (Equation 12) is unrealistic at large $|z|$ outside of the midplane, because (1) it does not include a contribution from the dark matter halo and (2) it assumes an infinite plane of stars. At large $|z|$, the dark matter halo will have more of an impact and the finite size of the galaxy will become important. Overall, these effects will decrease the vertical component of gravitational acceleration at large $|z|$, increasing the outflow rate we find in the simulations, helping to launch a large-scale cosmic-ray driven galactic wind.
Next, our radiative heating mechanism of a constant $\Gamma$ for all gas densities and vertical positions is unrealistic. The heating rate should fall off with vertical position $z$, allowing more radiative cooling to occur outside the midplane. This choice makes it harder for gas to cool at large $z$ where gas densities are also lower. However, choosing a constant $\Gamma$ makes the thermodynamic equilibrium curve (top panel of Figure 2) the same for the entire simulation volume, simplifying the analysis.
Also, we do not include cosmic-ray momentum boost of supernovae explosions (Diesing & Caprioli, 2018), or a local decrease in the cosmic-ray diffusion coefficient in star formation regions (Semenov et al., 2021). These additional complexities in cosmic-ray transport would primarily change cosmic-ray feedback in the midplane, but would have a smaller impact on larger length scales $(L\gtrsim 1\,\mathrm{kpc})$ and longer timescales $(t>1\,\mathrm{Myr})$ we consider.
Finally, our work does not complete the full stellar feedback loop to show how cosmic-ray injections impact the star formation rate. To complete that loop, the injection rate would at least need to be determined by star formation rates (e.g. like in the galaxy scale simulations of Chan et al. 2019 or other tallbox simulations like Rathjen et al. 2021; Kim et al. 2023). Because we do not complete the loop, the physical process we highlight of vertical magnetic field lines being cosmic ray escape highways is just a physical explanation of decreased $\gamma$-ray luminosity. The overall effect on star formation rate of magnetic field lines connecting the disk and halo is still unclear.
However, we suspect the true picture of cosmic ray feedback on star formation rates will require the increased magnetic field resolution. For example, reconsider Figure 6. In cases where the Parker instability is a dominant dynamical process, the decorrelation of cosmic rays and gas density could significantly impact start formation rates. However, that decorrelation was caused by field lines being vertical and not undergoing numerical reconnection. The vertical field lines also allowed cosmic rays to escape the midplane before they could interact with the cold gas, decreasing the $\gamma$-ray luminosity. So, the fundamental mechanism is the production of a vertical field, connecting our results to other setups which may not undergo the Parker instability but have a different way to produce vertical magnetic field.
@@ 5 Conclusion
The simulations we have run, CRInj and CRBkg, have illustrated the importance of cosmic-ray injections in driving outflows and structuring the ISM. Here, we list some key results from this study of cosmic-ray feedback:
@@
    The Parker instability drives overturning of gas which turns our initial hydrostatic equilibrium into a two-scale-height vertical gas density structure, in agreement with observations of the Milky Way. This overturning and saturation happens on a shorter timescale than previously predicted: $\sim 150\mathrm{Myr}$ instead of $\sim 500\mathrm{Myr}$ (Heintz et al., 2020)). This faster rate is the result of supernovae injection.
@@
    During the nonlinear saturation of the Parker instability, a more vertically oriented magnetic field forms just outside the midplane. This magnetic field structure allows for more efficient cosmic-ray escape alongside significant cosmic-ray feedback and a decrease in the calorimetric fraction $F_{\mathrm{cal}}$.
@@
    The same vertical magnetic field structure also allows for a cosmic ray pressure gradient driven outflow to survive for a long time period $(>100\,\mathrm{Myr})$. This outflow could form the base of a large scale galactic wind, stretching down to $|z|\sim 500\,\mathrm{pc}$.
  5. 5.
    The outflow is amplified in both the Parker instability regime and in the steady state that forms afterwards by the individual cosmic-ray injections included in the CRInj simulation.
  6. 6.
    We found that resolution was a significant factor in decreasing the calorimetric fraction $F_{\mathrm{cal}}$. Numerical reconnection of magnetic field lines disconnects the midplane from the high values of $|z|$, trapping cosmic rays in the midplane. Without high resolution of accurate subgrid modeling of vertical magnetic field, simulations will overpredict $\gamma$-ray luminosity of galaxies.
We would like to thank the anonymous reviewer, whose comments significantly improved the paper. In particular, their advice on supernova momentum feedback significantly improved our simulations. The authors would like to thank Sophie Aerdker, Bob Benjamin, Ryan Farber, Karol Fulat, Ka Wai Ho, David Hosking, Francisco Ley, Lukas Merten, Peng Oh, Nickolas Pingel, Mohan Richter-Addo, Mateusz Ruszkowski, Aaron Tran, Bindesh Tripathi, and Ka Ho Yuen for their insight and helpful conversations which improved this work. The authors greatly appreciate funding from funding from NASA FINESST grant No. 80NSSC22K1749 and NSF grant AST-2007323 which supported this work. Resources supporting this work were provided by the NASA High-End Computing (HEC) Program through the NASA Advanced Supercomputing (NAS) Division at Ames Research Center.